# Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding

### v1

@@ Abstract
Recent advances in vision-language models have significantly propelled video understanding. Existing datasets and tasks, however, have notable limitations. Most datasets are confined to short videos with limited events and narrow narratives. For example, datasets with instructional and egocentric videos often document the activities of one person in a single scene. Although some movie datasets offer richer content, they are often limited to short-term tasks, lack publicly available videos and frequently encounter data leakage given the use of movie forums and other resources in LLM training. To address the above limitations, we propose the Short Film Dataset (SFD) with 1,078 publicly available amateur movies, a wide variety of genres and minimal data leakage issues. SFD offers long-term story-oriented video tasks in the form of multiple-choice and open-ended question answering. Our extensive experiments emphasize the need for long-term reasoning to solve SFD tasks. Notably, we find strong signals in movie transcripts leading to the on-par performance of people and LLMs. We also show significantly lower performance of current models compared to people when using vision data alone.
@@ 1 Introduction
Recent advances in vision-language models [33, 31, 9, 10, 61, 60, 16, 15, 11, 68] show significant promise for enhancing machine perception. Despite their remarkable progress, current datasets are still constrained by some limitations. Most of them consist of short videos [20, 1, 13, 69, 72, 17, 55, 27, 32, 19, 65], often under one minute, and primarily target short-term tasks e.g., action recognition and video retrieval, often requiring only a few seconds of visual content for problem-solving [39, 28]. For long-form video understanding, the community has turned to three video categories. (i) Egocentric videos, an emerging domain featuring extended first-person sequences that capture a person’s continuous actions [18, 12, 24, 34, 39, 52, 51] or chronicle daily activities, offering tasks from object manipulation to social interactions [18]. Nevertheless, while such datasets excel in video duration, they lack narrative depth, focusing mainly on immediate visuals rather than rich storytelling; (ii) Instructional videos [41, 71], which depict a wide variety of human tasks and focus on the procedural aspect, are often brief and deficient in storytelling; (iii) Movies, unlike egocentric and instructional videos, provide complex plots and extended duration, making them an ideal testing ground for developing models that can comprehend long-form video narratives, see Figure 1.
@@
Several movie datasets have been proposed in the literature [48, 4, 56, 29, 53, 21, 64, 22, 59, 66, 49, 40, 44, 7, 6, 38] offering rich narrative-driven content and various tasks. These tasks include clip retrieval [4], video question answering [56, 29, 38], audio description [53, 21], semantic role labeling [49], scene segmentation [22, 44], and visual scene recognition [6]. However, current movie datasets suffer from three main limitations. (i) Accessibility: these datasets are not publicly available as they typically contain commercial copyrighted movies (see Figure 2). (ii) Clip duration: they often consist of incomplete and short clips [56, 29, 45, 64] (i.e., less than 4 minutes, see Figure 2), limiting their potential for long-form understanding with evolving storylines. Moreover, these datasets frequently suffer from ill-designed benchmark questions with insufficient narrative focus [54]. (iii) Data Leakage: as these datasets comprise well-known commercial movies, modern Large Language Models (LLMs) and Vision-Language Models (VLMs) have likely been exposed to some form of movie information (i.e., synopses, reviews, discussions, subtitles, blog posts). Figure 3 showcases this, where given only movie titles, modern LLMs can obtain high accuracy when answering questions from main movie datasets. This data leakage leads to biased benchmarks and ineffective training.
@@
To address the above limitations, we introduce the Short Film Dataset (SFD), a novel video benchmark comprising 1,078 short films, totalling over 243 hours of videos with an average duration of 13 minutes per film—significantly longer than existing datasets (see Figure 2). Unlike copyrighted films, SFD includes publicly accessible amateur films of diverse genres from YouTube. Although these films are shorter than traditional movies, they feature complex narratives that unfold through key sequences of events and multiple character interactions. Most importantly, with minimal exposure to LLMs, our dataset is less prone to data leakage issues plaguing other movie datasets (see Figure 3).
Aiming towards long-term story-level multimodal video understanding, we propose two question-answering tasks on SFD: (a) following recent trend [56, 45], we propose Multiple-Choice Question answering (MCQ); (b) We also introduce Open-Ended Question answering (OEQ), a more generative and unrestricted QA format that more realistically and objectively measures a model’s ability to comprehend long-form videos. All questions and answers are generated by LLMs from movie descriptions and undergo thorough manual curation, ensuring they accurately reflect the settings, characters, storylines, and themes of the movies.
To validate the advantages of SFD, we conduct extensive experiments and quantify data leakage, the benefits of long-term reasoning and comparison to human performance while using state-of-the-art vision and language models. Our findings are: (i) Unlike existing VideoQA datasets, SFD provides story-level tasks defined on public data with minimal data leakage (Section 3.1); (ii) our long-term reasoning experiments in Section 3.2 confirm the advantage of movie-level video understanding for solving SFD tasks; (iii) the evaluation in Section 3 reveals a large gap between recent VideoQA methods and human performance in vision-only and multimodal settings. SFD, hence, provides a unique benchmark for evaluating and advancing new methods for story-level long-term video understanding. Our dataset, code, and models are publicly available at https://shortfilmdataset.github.io.
2 Short Film Dataset
@@
Short films are motion pictures that typically last between 5 and 20 minutes. As illustrated in Figure 4, short films span a variety of styles (e.g., narrative fiction, documentary, animation, etc.) and genres (e.g., action, drama, comedy, horror, etc.). Filmmakers often use the short film format to explore new ideas, techniques, and storytelling styles. Short films, hence, have more of an experimental rather than commercial value and are often made publicly available. Samples from the dataset can be found in Appendix H.
The amount of publicly available short films has significantly increased over the last years (see Figure 6(c)). We take advantage of this development and create a Short Film Dataset (SFD) to foster research on story-level video understanding. In the remaining part of this section, we describe the procedure for collecting videos and corresponding metadata (Section 2.1) as well as generating question-answer pairs (Section 2.2). Section 2.3 presents the analysis and statistics of SFD. The overall pipeline for SFD creation is illustrated in Figure 5.
@@ 2.1 Data Collection
To build our dataset, we capitalize on the abundance of short films on video-sharing platforms. We specifically target the Omeleto YouTube channel111https://www.youtube.com/@Omeleto and its subchannels, including @OmeletoComedy and @OmeletoDrama, among others. These channels promote high-quality and award-winning short films, some of which have been recognized by Oscars and BAFTA awards. We download videos, subtitles, and associated metadata from YouTube using yt-dlp222https://github.com/yt-dlp/yt-dlp.
As speech transcripts were originally available only for a portion of the dataset, we have used WhisperX [3] to extract missing transcripts. The metadata accompanying each short film includes the movie title, a logline which is a concise one-sentence summary of the plot, and additional movie details such as genre, release year, region of origin, and language. Each film also comes with a detailed description that offers a deeper understanding of various elements such as a synopsis providing a brief narrative overview, the director’s inspirations, storytelling approach, actors’ performances, and other external information. To maintain clarity and objectivity in our dataset, we retain only the synopsis and exclude other subjective or ambiguous elements of the description.
2.2 Question and Answer Generation
To promote the development of long-form and story-oriented video understanding, we complement SFD with two Video Question Answering (VideoQA) tasks: Multiple-Choice Question answering (MCQ) and Open-Ended Question answering (OEQ). We first automatically generate question-answer pairs by prompting GPT-4 [42] with movie titles, loglines and synopses. We then manually curate questions and answers to ensure their relevance and correctness. We describe the generation process in more details below.
Automatic QA pair generation. For each film we obtain human-written information, i.e. movie titles, loglines and synopses from Omeleto YouTube channels. We then prompt GPT-4 [42] to propose relevant question-answer pairs from movie information, with dedicated prompts to avoid common mistakes. In particular, we emphasise direct and clear questions, with brief and specific answers to avoid speculation, repetition and ambiguity. Our example prompts are available in Appendix I. This stage results in 7,104 question-answer (QA) pairs that we use for the open-ended question-answering task.
Automatic Distractor Generation. To construct the MCQ task, we use GPT-4 [42] to complement each question with 4 distractors. The distractors are crafted to provide plausible yet incorrect answers that fit within the movie’s context. To this end, our prompts are designed with specific criteria: distractors should be syntactically similar to the correct answer while being semantically different. Our distractors also incorporate diverse misdirections relative to movies (character confusion, plot adjustments), consider the movie genre and plausible scenarios, and include accurate but irrelevant information from other parts of the synopsis. This stage results in 7,104 Multiple-Choice Questions (MCQs), with 5 options per question with one correct and four incorrect answers.
Manual QA Curation. To ensure the quality of questions and answers, we perform a thorough manual curation based on the following criteria. We remove questions that cannot be answered based on the synopsis, including factually incorrect, ambiguous, or subjective ones. Additionally, questions considered overly simplistic or answerable through external knowledge are excluded. This stage results in 4,885 MCQs from 1,078 movies.
2.3 Dataset Analysis and Statistics
Our SFD dataset contains a total of 1,078 unique films ranging in duration from 5 to 37 minutes and having an average length of 13 minutes (see Figure 6(a)). Overall, these movies amount to 243 hours and cover distinct genres, as illustrated in Figure 6(d). Each movie in the dataset is accompanied by a title, a logline which averages around 15 words, and a synopsis of about 97 words, equivalent to approximately 2 sentences, as shown in Figure 6(b). We use this metadata for question and answer generation as described in Section 2.2. The majority of films in SFD are in English and mainly originate from North America and Europe (see Figures 6(f) and 6(e)). Examples of the dataset’s diversity can be seen in Figure 4. Additional data preprocessing and statistics are available in Appendix C, including clips, captions, location tags, and face tracks.
In terms of video understanding tasks, SFD comprises 4,885 MCQs and OEQs, averaging 4.53 questions per film. Table 1 shows the comparison to both general video understanding (videoQA) and movie-oriented (movieQA) question-answering datasets. Notably, SFD exhibits the longest average video duration (821 seconds) among all datasets. Compared to videoQA datasets, SFD holds the properties of multimodality and long-term nature, similar to most movieQA datasets. Compared to other movieQA ones, SFD offers advantages in public availability of videos, data leakage (i.e., not exposed to LLMs), and story completeness (using full movies instead of trimmed clips, potentially impacting narrative coherence).
@@ 2.3 Dataset Analysis and Statistics
Dataset | Venue | Annotation | Avg. | #QA Pairs | Multimodal | Long-Term | Accessible | Unknown | Full |
Length (s) | To LLMs | Movies | |||||||
General VideoQA datasets | |||||||||
MSRVTT-QA [67] (test) | ACM 2017 | Auto | 15 | 72,820 | ✗ | ✗ | ✓ | ✓ | - |
MSVD-QA [8] (test) | ACM 2017 | Auto | 10 | 13,156 | ✗ | ✗ | ✓ | ✓ | - |
TGIF-QA [23] (test) | CVPR 2017 | Auto | 3 | 25,751 | ✗ | ✗ | ✓ | ✓ | - |
ActivityNet-QA [75] (test) | AAAI 2019 | Manual | 180 | 8,000 | ✗ | ✗ | ✓ | ✓ | - |
How2QA [50] (test) | EMNLP 2020 | Manual | 60 | 4,400 | ✗ | ✗ | ✓ | ✓ | - |
NeXT-QA [65] (test) | CVPR 2021 | Manual | 44 | 9,178 | ✗ | ✗ | ✓ | ✓ | - |
iVQA [35] | ICCV 2021 | Manual | 18 | 10,000 | ✗ | ✗ | ✓ | ✓ | - |
EgoSchema [39] | NeurIPS 2023 | Manual + Auto | 180 | 5,000 | ✗ | ✓ | ✓ | ✓ | - |
MovieQA datasets | |||||||||
MovieQA [56] (test) | CVPR 2016 | Manual | 203 | 6,462 | ✓ | ✓ | ✗ | ✗ | ✗ |
TVQA [29] (test) | EMNLP 2018 | Manual | 76 | 15,253 | ✓ | ✗ | ✗ | ✓ | ✗ |
LVU [64] (test) | CVPR 2021 | Manual | 220 | 1,223 | ✓ | ✓ | ✓ | ✗ | ✗ |
MovieChat [54] (test) | CVPR 2024 | Manual | 459 | 2,417 | ✓ | ✓ | ✗ | ✓ | ✗ |
CinePile [45] (test) | arXiv 2024 | Manual + Auto | 160 | 4,940 | ✓ | ✓ | ✓ | ✓ | ✗ |
SFD (Ours) | Manual + Auto | 821 | 4,885 | ✓ | ✓ | ✓ | ✓ | ✓ |
Question Analysis.
The generated question-answer pairs are long and diverse: the median answer length is 10 words and the vocabulary contains 8,928 different words. Furthermore, to ensure that the generated QAs cover various aspects of movies, we categorize them into four types (see Table 2):
1. Setting-related questions focus on the location and time of the movie. They may pertain to geographical locations (e.g., a specific city or country), a type of place (e.g., home, office), a historical period (e.g., a specific era), or a time of day (e.g., morning, night).
2. Character-related questions focus on the individuals portrayed in the movie. They may concern personal information (age, gender, profession), traits, motivations, and relationships (siblings, friends). Note that character development, as part of the narrative, is excluded from this category.
3. Story-related questions focus on the key sequence of events and interactions that compose the plot and narrative arc of the movie.
4. Theme-related questions focus on the underlying message or central idea that the movie explores and conveys to its audience.
Question Type | Questions | Answers |
Setting | Where do John and Emma have their conversation? | On the bus. |
Which city is the backdrop? | Liverpool. | |
What decade does the movie take place in? | 1990s. | |
Character | Who is the main character? | Josh. |
What is the nationality of the soldier who becomes a prisoner? | German. | |
How is Mr. Jones described? | As stern and harsh. | |
Story | Can you name a specific tactic Tommy uses against Tiny? | He steals Tiny’s food at dinner and distracts her at a school-wide relay race. |
What is the imminent threat to Greg’s residence? | The house is in foreclosure and the bank is about to repossess it. | |
What triggers Carla to believe her husband Frank will surprise her for their anniversary? | A bouquet left on her doorstep. | |
Theme | What is a major theme? | The theme revolves around reconnecting and the complexities of honesty between exes. |
What main theme is explored through Sarah’s experience? | The theme of romantic expectations versus present realities. | |

### v2

# Long Story Short: Story-level Video Understanding from 20K Short Films

@@ Abstract
Recent developments in vision-language models have significantly advanced video understanding. Existing datasets and tasks, however, have notable limitations. Most datasets are confined to short videos with limited events and narrow narratives. For example, datasets with instructional and egocentric videos often depict activities of one person in a single scene. Although existing movie datasets offer richer content, they are often limited to short-term tasks, lack publicly available videos, and frequently encounter data leakage issues given the use of subtitles and other information about commercial movies during LLM pretraining. To address the above limitations, we propose Short-Films 20K (SF20K), the largest publicly available movie dataset. SF20K is composed of 20,143 amateur films and offers long-term video tasks in the form of multiple-choice and open-ended question answering. Our extensive analysis of SF20K reveals minimal data leakage, emphasizes the need for long-term reasoning, and demonstrates the strong performance of recent VLMs. Finally, we show that instruction tuning on the SF20K-Train set substantially improves model performance, paving the way for future progress in long-term video understanding.
@@ 1 Introduction
Recent advances in vision-language models [35, 33, 10, 11, 68, 65, 19, 18, 12, 75] show significant promise for enhancing machine perception. Despite their remarkable progress, current datasets are still constrained by some limitations. Most of them consist of short videos [23, 1, 15, 76, 78, 20, 60, 29, 34, 22, 71], typically under one minute, and primarily target short-term tasks such as action recognition and video retrieval, which often require only a few seconds of visual content to solve [41, 30]. Long-form video understanding, on the other hand, has been mainly studied in three domains: (i) Egocentric videos, featuring extended first-person sequences with continuous actions [21, 13, 27, 37, 41, 57, 56] and daily activities ranging from object manipulation to social interactions [21]. Although these datasets excel in video duration, they lack narrative depth, focusing on immediate visuals rather than rich storytelling; (ii) Instructional videos [44, 77], which depict a variety of human tasks and focus on the procedural aspect, are often brief and deficient in storytelling; (iii) Movies, unlike egocentric and instructional videos, provide complex plots and extended durations, making them well suited for developing models capable of comprehending long-form narratives (see Figure 1).
@@
Several movie datasets have been proposed in the literature [53, 5, 61, 31, 58, 24, 70, 25, 64, 73, 54, 42, 48, 8, 7, 40, 59, 50, 72] offering rich, narrative-driven content and supporting various tasks. These tasks include clip retrieval [5], video question answering [61, 31, 40, 59, 50], audio description [58, 24, 72], semantic role labeling [54], scene segmentation [25, 48], and visual scene recognition [7]. However, current movie datasets suffer from three main limitations: (i) Accessibility: access to commercial movies is subject to copyright restrictions (see Figure 2); (ii) Clip duration: they often consist of incomplete and short clips [61, 31, 49, 70], often less than four minutes in length (see Figure 2), limiting their potential for long-form understanding of evolving storylines. Moreover, these datasets frequently suffer from ill-designed benchmark questions with insufficient narrative focus [59]; (iii) Data leakage: as these datasets feature well-known commercial movies, modern Large Language Models (LLMs) and Vision-Language Models (VLMs) have likely been exposed to related content, such as synopses, reviews, discussions, subtitles, and blog posts. Figure 3 showcases this issue, showing that modern LLMs can achieve high accuracy on questions from major movie datasets using only movie titles. This leakage leads to biased benchmarks and ineffective training.
@@
To address the limitations of existing movie datasets, we introduce Short-Films 20K (SF20K), a novel video dataset comprising 20,143 short films totaling almost 3,582 hours of video, with an average duration of 11 minutes per film — significantly longer than existing datasets (see Figure 2). Unlike copyrighted films, SF20K consists of publicly accessible amateur films across diverse genres sourced from YouTube and Vimeo. While shorter than traditional movies, these films feature complex narratives that unfold through key sequences of events and multiple character interactions. Crucially, with minimal exposure to LLMs, SF20K is less prone to data leakage issues plaguing other movie datasets (see Figure 3).
Aiming towards long-term story-level video understanding, we propose two question-answering tasks on SF20K: (a) Multiple-Choice Question Answering (MCQA), following recent trends [61, 49]; and (b) Open-Ended Question Answering (OEQA), a less restrictive QA format requiring free-form answers. The questions and answers are initially generated using LLMs based on movie descriptions. For the test set, all QAs undergo meticulous manual curation to ensure they accurately reflect the settings, characters, storylines, and themes of the films. To demonstrate the benefits of SF20K, we conduct extensive experiments including an analysis of data leakage, the impact of long context windows, model performance relative to human benchmarks using state-of-the-art vision-language models, and the effect of instruction tuning on the training set.
Our contributions are threefold: (1) We propose SF20K, the largest publicly accessible movie dataset, comprising 20,143 short films and offering two story-level videoQA tasks. (2) Through extensive experiments, we show: (i) Unlike other movie datasets, SF20K exhibits minimal data leakage to modern LLMs; (ii) Recent video models show significantly lower performance compared to humans, particularly in the multimodal setting; (iii) A long temporal window is required to solve the QA tasks, highlighting the need for models capable of handling information at the movie level. (3) We demonstrate that large-scale instruction tuning on SF20K-Train helps model performance. We argue that SF20K provides a unique dataset for advancing long-term, story-level video understanding. The dataset, code, and models are publicly available at https://ridouaneg.github.io/sf20k.html.
2 Related Work
Video QA benchmarks. Questions in such benchmarks can be framed to assess the reasoning, memory, and comprehension skills of both humans and models. Hence, many datasets have been made available through the years, covering visual descriptions [76, 9], temporal action reasoning [71], compositional reasoning [22], social intelligence [82], instructional videos [77], egocentric videos [41, 27, 37], movies [61, 31, 40], among others [26, 74, 26, 34, 81, 69]. However, these datasets feature short videos and typically require reasoning on only a few frames to solve the task at hand, as shown by temporal certificates in [41].
Long-form video understanding. As new long-context models emerge [51], we need more complex benchmarks. Long-form video understanding evaluates long-term reasoning capabilities of those models. Some datasets have been introduced in recent years, for egocentric videos [41], movies [70, 58, 24, 61, 50], and general videos [17, 66]. For instance, [41] proposes multiple-choice questions on 3-minute videos that require on average 100 seconds of viewing to be answered correctly. [70] includes long movie clips, however the subsequent tasks are relatively limited, focusing on less relevant aspects such as ’like’ ratio and view count prediction. [58, 24] introduce an audio description task, which is currently tackled at the clip level but could be extended over the whole movie. Recently, [17, 66] introduced longer videos that present significant challenges even for the most advanced vision-language models. [61, 59] propose movie question-answering datasets focused on movie plots, settings and characters, though access to the data is limited. Finally, closer to our work, [50] offers a dataset of question-answer pairs derived from commercial movie clips, providing both training and test sets; however, along with other movie datasets, it may suffer from data leakage issues.
3 SF20K: A Dataset of 20K Short Films
@@
Short films are motion pictures that typically last between 5 and 40 minutes. As illustrated in Figure 4, short films span a variety of styles (e.g., narrative fiction, documentary, animation) and genres (e.g., drama, comedy, horror). Filmmakers often use the short film format to explore new ideas, techniques, and storytelling styles. Short films, hence, have more of an experimental rather than commercial value and are often made publicly available.
The amount of publicly available short films has significantly increased over the last few years (see Figure 5(b)). We take advantage of this development and create SF20K, a dataset of 20,143 short movies collected from online platforms, totaling 3,582 hours of video content, with each video lasting on average 11 minutes. Our procedure for collecting videos, metadata, and annotations for SF20K is described in Section 3.1.
To promote the development of long-form, story-oriented video understanding, we introduce two video QA tasks: Multiple-Choice Question Answering (MCQA) and Open-Ended Question Answering (OEQA). The QA generation process — consisting of synopsis preparation, automatic question generation, and manual curation — is presented in Section 3.2. Section 3.3 presents detailed statistics about SF20K, including annotations, tasks, and comparison to other datasets. Section 3.4 presents a specific analysis of SF20K-Test, including question types and subsets.
@@ 3.1 Data Collection
Source and filtering. To build SF20K, we capitalize on the abundance of short films on video-sharing platforms. We target YouTube and Vimeo channels hosting high-quality and award-winning short films, some of which have been recognized by Oscars and BAFTA awards 111Among others, we use the YouTube channels Omeleto https://www.youtube.com/@Omeleto, Film Shortage https://www.youtube.com/@FilmShortage, and the Shortverse library https://www.shortverse.com. Videos, subtitles, and metadata are downloaded using yt-dlp 222https://github.com/yt-dlp/. We filter out videos less than three months old (before July 2024), and rely on platform safety filters to exclude sensitive, age-restricted, and private content.
Annotation extraction. Since speech transcripts are available for only a portion of the dataset, we use Whisper Large-v3-Turbo [47] to generate missing subtitles. The metadata accompanying each short film includes the movie title, logline (a concise one-sentence plot summary), genre, release year, region of origin, and language. Additionally, we extract shot boundaries using PySceneDetect333https://github.com/Breakthrough/PySceneDetect, face tracks using DeepFace444https://github.com/serengil/deepface, and frame captions using Llama-3.2-Vision-11B [43]. More details can be found in Appendix B.
3.2 Scaling Up Instruction Tuning
SF20K includes two tasks: Multiple-Choice Question Answering (MCQA) and Open-Ended Question Answering (OEQA). For both tasks, we use movie titles, loglines, and synopses — manually curated for the test set and automatically extracted for the training set (Section 3.2.1). Given this data, we automatically generate question-answer pairs by prompting GPT-4 [45] (Section 3.2.2). For SF20K-Test, we further refine the QA pairs through manual curation to ensure their relevance and correctness, and we augment this test set with additional manually created QA pairs (Section 3.2.3). We describe the generation process in detail below. The prompts used for question generation are available in Appendix H.
3.2.1 Synopsis Preparation
Human-written synopses for SF20K-Test. Among the 20K short films in SF20K, 1,072 come with detailed, human-written synopses. We use these high-quality synopses in the QA generation pipeline for the test set.
Synthetic synopses for SF20K-Train. For the training set, we generate synthetic synopses by prompting GPT-4 [45] using subtitles. This process captures key story-level details, producing one synopsis per movie.
3.2.2 Automatic Question Generation
The automatic QA generation process is the same for both the train and test sets. See Appendix H for the exact prompts.
QA pair generation. GPT-4 [45] is used to generate question-answer pairs based on movie titles, loglines, and synopses. Prompts are carefully designed to obtain concise, specific, and unambiguous questions, with brief answers.
Distractor generation. To construct the MCQA task, we use GPT-4 [45] to complement each question-answer pair with four distractors. Distractors are crafted to be plausible yet incorrect, using syntactic similarity to the correct answer while maintaining semantic distinctions. Diverse misdirections, such as character confusion and irrelevant but accurate details, are included.
3.2.3 Manual Curation
SF20K-Train. The training set relies entirely on automatically generated QA pairs without manual curation.
SF20K-Test. The test set undergoes manual curation to ensure high quality. Ambiguous, subjective, or overly simplistic questions are removed.
@@ 3.2.3 Manual Curation
| Dataset | Venue | Annotation | Average | Total | #QA Pairs | Multimodal | Long Term | Accessible | Unknown | Full | |
| Duration (s) | Hours | Train | Test | To LLMs | Movies | ||||||
| VideoQA datasets | |||||||||||
| MSRVTT-QA [74] | ACM 2017 | Auto | 15 | 41 | 170,860 | 72,820 | ✗ | ✗ | ✓ | ✓ | - |
| MSVD-QA [9] | ACM 2017 | Auto | 10 | 5 | 37,344 | 13,156 | ✗ | ✗ | ✓ | ✓ | - |
| TGIF-QA [26] | CVPR 2017 | Auto | 3 | 47 | 139,414 | 25,751 | ✗ | ✗ | ✓ | ✓ | - |
| ActivityNet-QA [81] | AAAI 2019 | Manual | 180 | 290 | 32,000 | 8,000 | ✗ | ✗ | ✓ | ✓ | - |
| How2QA [55] | EMNLP 2020 | Manual | 60 | 366 | 35,205 | 4,400 | ✗ | ✗ | ✓ | ✓ | - |
| NeXT-QA [71] | CVPR 2021 | Manual | 44 | 66 | 71,655 | 17,742 | ✗ | ✗ | ✓ | ✓ | - |
| iVQA [38] | ICCV 2021 | Manual | 18 | 50 | - | 10,000 | ✗ | ✗ | ✓ | ✓ | - |
| EgoSchema [41] | NeurIPS 2023 | Auto + Manual | 180 | 250 | - | 5,000 | ✗ | ✓ | ✓ | ✓ | - |
| MovieQA datasets | |||||||||||
| MovieQA [61] | CVPR 2016 | Manual | 203 | 380 | 4,318 | 1,258 | ✓ | ✓ | ✗ | ✗ | ✗ |
| TVQA [31] | EMNLP 2018 | Manual | 76 | 460 | 122,036 | 15,253 | ✓ | ✗ | ✗ | ✓ | ✗ |
| MovieChat [59] | CVPR 2024 | Manual | 459 | 180 | - | 14,000 | ✓ | ✓ | ✗ | ✓ | ✗ |
| CinePile [49] | arXiv 2024 | Auto + Manual | 160 | 417 | 298,887 | 4,941 | ✓ | ✓ | ✓ | ✓ | ✗ |
| SF20K (Ours) | - | Auto + Manual | 640 | 3,582 | 191,007 | 4,885 | ✓ | ✓ | ✓ | ✓ | ✓ |
3.3 Dataset Statistics
SF20K contains 20,143 short films ranging from 5 to 37 minutes, with an average length of 11 minutes per video (see Figure 5(a)), and totaling 3,582 hours. The dataset covers multiple genres, as illustrated in Figure 5. Examples of the dataset’s diversity can be seen in Figure 4. Most of the films are in English and originate mainly from North America and Europe (see Figure 5(d,e)). More samples can be seen in Appendix I.
Annotations. SF20K includes 2.4M shots, with an average shot length of 5.52 seconds. Each shot has a 24-word caption describing settings and characters. Additional annotations include 1.1M subtitles and 1.08M face tracks.
Tasks. SF20K-Train contains 191,007 questions, each suitable for both multiple-choice and open-ended formats, averaging 9.21 questions per movie. SF20K-Test includes 4,885 curated QA pairs, averaging 4.53 questions per movie.
Comparison to other datasets. Table 1 compares SF20K to existing datasets. Notably, SF20K exhibits the longest average duration (640 seconds). Compared to video QA benchmarks, SF20K-Test distinguishes itself by its multimodal and long-term focus. Compared to movie QA benchmarks, SF20K-Test offers advantages in public availability of videos, limited data leakage (unlikely exposure to LLMs), and story completeness, as it uses full movies rather than trimmed clips, potentially impacting narrative coherence.
3.4 Test Set Analysis
This section analyzes SF20K-Test, focusing on question types and subset descriptions.
Question types. SF20K-Test contains four question types:
Setting-related questions address location and time of the movie, including geographical places, types of locations (e.g. home, office), historical periods, and times of day.
Character-related questions focus on personal details (age, profession), traits, and relationships of the main characters.
Story-related questions cover key events driving the plot.
Theme-related questions explore the movie’s central idea.
Test subsets. SF20K-Test comes with two specific subsets:
1. SF20K-Test-Silent. This subset contains silent films with minimal or no dialogue, which accounts for around 10% of SF20K-Test. It consists of 90 movies (20.5 hours) and 419 questions. This subset is intended to challenge vision-language models that heavily rely on language information.
2. SF20K-Test-Expert. We ask human annotators to watch movies and craft questions, capturing complex elements like implicit information, key visual details essential to the story, and plot twists. This subset contains 50 movies (11.4 hours) and 570 questions, averaging 11.4 questions per movie (compared to 4.55 in SF20K-Test overall) and is designed exclusively for open-ended question answering.